# Rings of non-commutative functions and their fields of fractions

### v1

@@ 2.2 Bounded NC functions
The non-commutative Hardy algebra, $\mathbb{H}^{\infty}_{d}$, is the algebra of all uniformly bounded NC functions in the NC unit row-ball, $\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$. This algebra can be identified with the unital weak operator topology (WOT)–closed algebra generated by the $L_{1},\ldots,L_{d}$ [13]. Moreover, $\mathbb{H}^{\infty}_{d}$ can be identified with the algebra of left multipliers of the NC-RKHS $\mathbb{H}^{2}_{d}$ [41, 43].
@@
This simple fact can be used to prove the following lemma, [24, Lemma 4.5].
@@
$=$ $\sup\left\{\left.\langle{K\{Z,y,v\}},{f}\rangle\right|\ \|v\|,\,\|y\|=1\right\}$
@@
For any $f\in\mathbb{H}^{\infty}_{d}$ we define $f(L):=M^{L}_{f}\in\mathscr{B}(\mathbb{H}^{2}_{d})$ as the bounded linear operator that acts as left multiplication by $f$. We note that for every $0<r<1$ the operator $rL$ is a strict row contraction. Hence, the bounded linear operator, $f(rL)$, is well-defined. The operator $f(rL)$ corresponds to the restriction of $f$ to $r\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$. More generally, this identification extends to $f\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{n\times k}$, for all sizes of matrices.
@@
$F(rL)$ will be surjective if and only if $F(rL)^{*}$ is bounded below. We first claim that $F(rL)^{*}$ is bounded below if $F(Z)^{*}$ is uniformly bounded below for all $Z\in r\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$.
@@
Suppose that $F(rL)^{*}$ is not bounded below. Then, given any $\epsilon>0$ there exists an $x\in\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{k}$ so that $\|F(rL)^{*}x\|<\frac{\epsilon}{2}\|x\|$. Since $F(rL)\in\mathbb{A}_{d}\otimes\mathbb{C}^{n\times k}$, it can be approximated in operator norm by its Cesàro sums. Hence there exists an $N\in\mathbb{N}$ so that $\ell>N$ implies that $\|F_{\ell}(rL)^{*}-F(rL)^{*}\|<\frac{\epsilon}{2}$, where $F_{\ell}(rL)$ denotes the $\ell$-th Cesàro sum. Finally, we can approximate the “infinite–dimensional point”, $rL$, by finite–dimensional compressions, $Z^{(j)}\in r\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, of size $j$, by compressing $rL$ by the projection, $P_{j}$ onto the first $j$ standard basis vectors of $\mathbb{H}^{2}_{d}$. Then, since $F_{\ell}\in\mathbb{C}\langle\mathfrak{z}\rangle\otimes\mathbb{C}^{k\times n}$, $F_{\ell}(Z^{(j)})^{*}$ converges, in the strong operator topology, to $F_{\ell}(rL)^{*}$. Hence, there is a $J\in\mathbb{N}$ so that $j>J$ implies that
@@
The last part of the lemma follows from Popescu’s corona theorem [36, Corollary 3.2]. Since $T(rL)$ is surjective, $T(rL)^{*}$ is bounded below, and thus the corona theorem guarantees the existence of a right inverse. ∎
We note that if $f$ is an NC entire function then the operator $f(rL)$ makes sense for every $r>0$. Moreover, since the power series of $f$ converges uniformly on any ball, we have $f(rL)\in\mathbb{A}_{d}$. Here, $\mathbb{A}_{d}$, the NC or free disk algebra of Popescu is the norm closure of the free algebra in $\mathbb{H}^{\infty}_{d}$. Equivalently, $\mathbb{A}_{d}$ is completely isometrically isomorphic to the unital norm closed algebra generated by the left creation operators in $\mathscr{B}(\mathbb{H}^{2}_{d})$. We will typically view elements of $\mathbb{A}_{d}$ and $\mathbb{H}^{\infty}_{d}$ as left multiplication operators on $\mathbb{H}^{2}_{d}$ and given any $F\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{m\times n}$ we will sometimes write $F(L)$ in place of $F$, to denote that $F$ acts as left multiplication by $F$. Indeed, evaluating the free formal power series of $F$ at $L=(L_{1},\cdots,L_{d})$ yields the operator of left multiplication by $F$. (If $F\in\mathbb{A}_{d}\otimes\mathbb{C}^{m\times n}$ the partial Cesàro sums of this power series converge in operator-norm while if $F\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{m\times n}$ these Cesàro sums and their adjoints converge in the strong operator topology.)
@@
Let $f_{1},\ldots,f_{n}\in\mathscr{O}_{d}(R)$, $R\in(0,+\infty]$. Let $F=(f_{1},\cdots,f_{n})$. For every $0<r<R$, let $F(rL)=V_{r}(L)T_{r}(L)$ be the inner-outer decomposition. Then, there exists $0<\rho<R$ and $\leq n, kk\in\mathbb{N}$, such that for every $r>\rho$, the length of $V_{r}$ is $k$. Moreover, if $\rho<r<s<R$, then $V_{s}(\tfrac{r}{s}L)$ is injective and $\mathrm{Ker}\,F(rL)=\mathrm{Ker}\,T_{s}(\tfrac{r}{s}L)$.
@@
Since the dimension of the wandering subspace is, at most, the number of generators, we obtain that $w(r)\leq w(s)$. In other words, the function $w(t)$ is monotonically increasing, bounded, and integer–valued. Therefore, there exists a $\rho>0$, such that for all $0<\rho<r<R$, $w(r)$ is constant.
Now for $r>\rho$, we apply the inner-outer decomposition $V_{s}(\tfrac{r}{s}L)=VH_{s,r}$, where $H_{s,r}$ is outer and square. Observe that any square outer is necessarily injective. Indeed, at every point $Z\in\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, we have that $H_{s,r}(Z)$ is surjective and, therefore, invertible. We conclude that any vector in the kernel of $H_{s,r}$ must evaluate to 0 on $\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ and, thus, is zero. Therefore, $V_{s}(\tfrac{r}{s}L)$ is injective.
The last statement then follows as if $V_{s}(\tfrac{r}{s}L)$ is injective, $T_{s}(\tfrac{r}{s}L)$ is surjective and $F(rL)=V_{s}(\tfrac{r}{s}L)T_{s}(\tfrac{r}{s}L)$ is injective, we must also have that $T_{s}(\tfrac{r}{s}L)$ is injective. By the open mapping theorem, $T_{s}(\tfrac{r}{s}L)$ is then invertible. Any invertible $A(L)\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{m\times n}$ must be square, i.e. $m=n$. Indeed, $A(L):\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{n}\rightarrow\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{m}$ is surjective. Hence $\{A(L)1\otimes e_{k}\}_{k=1}^{n}$ is a $R\otimes I_{m}$-cyclic set for $\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{m}$ and hence $m\leq n$. Conversely, $A(L)^{-1}:\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{m}\rightarrow\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{n}$ is also surjective, so $m=n$. ∎
@@
[24, Appendix] Let $E\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{n\times n}$ be idempotent. Then there exists an $S\in\operatorname{GL}_{n}(\mathbb{H}^{\infty}_{d})$ and a $1\leq k\leq n$ so that
@@
This lemma was established in [24, Appendix], we include the proof for completeness.
@@
Write $S(L):=(V(L),W(L))$. Since $E(L)^{2}=E(L)$, we get that $V(L)(I_{\mathbb{H}^{2}_{d}}\otimes I_{k}-F(L)V(L))F(L)=0$. Since $V(L)$ is injective and $F(L)$ has dense range, we get that $F(L)V(L)=I_{\mathbb{H}^{2}_{d}}\otimes I_{k}$. Similarly, $G(L)W(L)=I_{\mathbb{H}^{2}_{d}}\otimes I_{j}$. Moreover, since E(L)(IE(L))=0=(IE(L))E(L)E(L)(I-E(L))=0=(I-E(L))E(L), we get that $F(L)W(L)=G(L)V(L)=0$. Hence, $S(L)$ is invertible with inverse $S(L)^{-1}=\left(\begin{smallmatrix}F(L)\\ G(L)\end{smallmatrix}\right)$. In particular, both are square, i.e., $j+k=n$.
@@
One of our main results, Theorem 4.3, proves that each $\mathscr{O}_{d}(R)$, $R\in(0,+\infty]$, is a semifir, and hence is projective–free. That is, a ring is projective–free if every finitely generated projective module over the ring is a free module. This implies, in particular, that any idempotent $E\in\mathscr{O}_{d}(R)^{n\times n}$ has such a factorization. Indeed, given such an $E$, we have that $\mathrm{Ran}\,E$ and $\mathrm{Ran}\,I-E$ are both finitely–generated projective right $\mathscr{O}_{d}(R)$-modules contained in $\mathscr{O}_{d}(R)^{n}$ obeying $\mathscr{O}_{d}(R)^{n}=\mathrm{Ran}\,E+\mathrm{Ran}\,I-E$ and $\mathrm{Ran}\,E\cap\mathrm{Ran}\,I-E=\{0\}$. Since $\mathrm{Ran}\,E$, $\mathrm{Ran}\,I-E$ are fintely–generated and $\mathscr{O}_{d}(R)$ is semifir, these are both free right $\mathscr{O}_{d}(R)$-modules of finite rank. Choosing a basis $\{f_{1},\cdots,f_{k}\}\subseteq\mathscr{O}_{d}(R)$ for $\mathrm{Ran}\,E$ and a basis $\{f_{k+1},\cdots,f_{n}\}\subseteq\mathscr{O}_{d}(R)$ for $\mathrm{Ran}\,I-E$, it follows that $\{f_{i}\}_{1}^{n}$ is a basis for $\mathscr{O}_{d}(R)^{n}$. Defining the matrix $S\in\mathscr{O}_{d}(R)^{n\times n}$ by
@@
Let $S\in\operatorname{GL}_{n}(\mathbb{H}^{\infty}_{d})$. Let $\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{n}=\mathcal{H}+\mathcal{K}$ be a decomposition into a non-orthogonal closed direct sum of right invariant subspaces. Then, $\dim\mathscr{W}_{\mathcal{H}}+\dim\mathscr{W}_{\mathcal{K}}=n$, $\dim\mathscr{W}_{S\mathcal{H}}=\dim\mathscr{W}_{\mathcal{H}}$, and $\dim\mathscr{W}_{S\mathcal{K}}=\dim\mathscr{W}_{\mathcal{K}}$.
@@
The first claim is proved in [24, Appendix] and in Lemma 2.6 above. For the second claim, we note that $S\mathscr{W}_{\mathcal{H}}$ generates $S\mathcal{H}$ by right shifts. Therefore, $\dim\mathscr{W}_{S\mathcal{H}}\leq\dim\mathscr{W}_{\mathcal{H}}$. Now applying this inequality to $S^{-1}$, we get the equality. Similarly, for $\mathcal{K}$. ∎
The following lemma is a variation of [24, Lemma 5.2].
@@
Let $0<r<s<R$. Then the map $\Phi_{r/s}$ will map $F(sL)$ to $F(rL)$. Therefore, for every $h\in\mathrm{Ker}\,F(sL)$, we have that $\Phi_{r/s}(h)\in\mathrm{Ker}\,F(rL)$. Hence, $\mathcal{H}:=\overline{\Phi_{r/s}(\mathrm{Ker}\,F(sL))}$ is a closed right-invariant subspace of $\mathrm{Ker}\,F(rL)$. In particular, $W:=\Phi_{r/s}(\mathscr{W}_{\mathrm{Ker}\,F(sL)})$ is a generating set for $\mathcal{H}$. Let $P$ be the projection onto $\mathscr{W}_{\mathcal{H}}$. Let $u\in\mathscr{W}_{\mathcal{H}}\ominus PW$. Fix an arbitrary $w\in W$. Then,
@@
Since $\mathcal{H}\ominus\mathscr{W}_{\mathcal{H}}$ is $R$-invariant, we get that the second summand on the left is 0. Since $R^{\alpha*}u=0$ for every non-empty $\alpha$ and $u\perp w$ by assumption, we get that the first summand is 0. Since $W$ is a generating set, we get that $u=0$. Therefore, $\mathscr{W}_{\mathcal{H}}=\overline{PW}$. ∎
@@
In [24, Appendix], it is proved that $\mathbb{H}^{\infty}_{d}$ is a projective free ring. However, it is also claimed in [24, Appendix] that $\mathbb{H}^{\infty}_{d}$ is a semifir. This fails even for $d=1$, as can be seen in [52, Beispiel 2]. To be more precise, if we take $f(z)=e^{\frac{z+1}{z-1}}$ and g(z)=1zg(z)=1-z, then by the corona theorem $f$ and $g$ do not generate the trivial ideal. Moreover, the ideal generated by the two is not principal. The reason is that if $h$ divides both $f$ and $g$, then we can write $h=h_{i}h_{o}$, where $h_{i}$ is inner and $h_{o}$ is outer. Since $f=ha$, then $h_{o}$ must be invertible since $f$ has no outer part. On the other hand, since $g=hb$, $h_{i}=1$ since $g$ has no inner part. Hence, $h$ is a unit, but this contradicts the fact that the ideal generated by $f$ and $g$ is not trivial.
As we will see, the obstruction that prevents $\mathbb{H}^{\infty}_{d}$ from being a semifir is that it is generally not possible to “trivialize” row outers that have dense range in the sense of [10, Section 2.3 and Theorem 2.3.1]. In this paper, we will see that for outers this problem can be overcome by increasing the radii of the balls. Namely, any such row outer on a row-ball of a given radius can always be trivialized on any ball of strictly smaller radius. Even though $\mathbb{H}^{\infty}_{d}$ is not a semifir, it would be interesting to determine whether $\mathbb{H}^{\infty}_{d}$ is a Sylvester domain, which is a larger class of rings than semifirs that also have universal skew fields of fractions [10, Section 5.5].

### v2

@@ 2.2 Bounded NC functions
The non-commutative Hardy algebra, $\mathbb{H}^{\infty}_{d}$, is the algebra of all uniformly bounded NC functions in the NC unit row-ball, $\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$. This algebra can be identified with the unital weak operator topology (WOT)–closed algebra generated by $L_{1},\cdots,L_{d}$ [13]. Moreover, $\mathbb{H}^{\infty}_{d}$ can be identified with the algebra of left multipliers of $\mathbb{H}^{2}_{d}$ [43, 45].
@@
This simple fact can be used to prove the following lemma, [25, Lemma 4.5].
@@
$=$ $\sup\left\{\left.|\langle{K\{Z,y,v\}},{f}\rangle|\ \right|\ \|v\|,\,\|y\|=1\right\}$
@@
For any $f\in\mathbb{H}^{\infty}_{d}$ we define $f(L):=M^{L}_{f}\in\mathscr{B}(\mathbb{H}^{2}_{d})$ as the bounded linear operator that acts as left multiplication by $f$. We note that for every $0<r<1$ the operator $rL$ is a strict row contraction. Hence, the bounded linear operator, $f(rL)$, is well-defined. The operator $f(rL)$ corresponds to the restriction of $f$ to $r\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$. More generally, this identification extends to $f\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{n\times k}$, for all sizes of matrices.
@@
Since $F$ is outer, $F(L)$ has dense range, by definition, and it is clear that $F(rL)$ will then also have dense range for any $0<r<1$. Moreover, $F(rL)$ will be surjective if and only if $F(rL)^{*}$ is bounded below. We first claim that $F(rL)^{*}$ is bounded below if $F(Z)^{*}$ is uniformly bounded below for all $Z\in r\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$.
@@
Suppose that $F(rL)^{*}$ is not bounded below. Then, given any $\epsilon>0$ there exists an $x\in\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{k}$ so that $\|F(rL)^{*}x\|<\frac{\epsilon}{2}\|x\|$. Since $F(rL)\in\mathbb{A}_{d}\otimes\mathbb{C}^{n\times k}$, it can be approximated in operator norm by matrices of free polynomials. In particular, it is the operator–norm limit of its Cesàro partial sums. This claim can be proven from the properties of the Cesàro summation map [13], the fact that for a function $f\in\mathbb{A}_{d}$, $f(tL)$ converges to $f$ in norm as $t$ increases to $1$ [40], a straightforward estimate that shows that the free polynomial partial sums of $f(tL)$ converge in norm to $f(tL)$ for any $t\in(0,1)$, and a standard $\varepsilon/3$-argument. Hence, there exists an $N\in\mathbb{N}$ so that $\ell>N$ implies that $\|F_{\ell}(rL)^{*}-F(rL)^{*}\|<\frac{\epsilon}{2}$, where $F_{\ell}(rL)$ denotes the $\ell$-th Cesàro sum. Finally, we can approximate the “infinite–dimensional point”, $rL$, by finite–dimensional compressions, $Z^{(j)}\in r\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, of size $j$, by compressing $rL$ by the projection, $P_{j}$ onto the first $j$ standard basis vectors of $\mathbb{H}^{2}_{d}$. Then, since $F_{\ell}\in\mathbb{C}\langle\mathfrak{z}\rangle\otimes\mathbb{C}^{k\times n}$, $F_{\ell}(Z^{(j)})^{*}$ converges, in the strong operator topology, to $F_{\ell}(rL)^{*}$. Hence, there is a $J\in\mathbb{N}$ so that $j>J$ implies that
@@
The last part of the lemma follows from Popescu’s corona theorem [37, Corollary 3.2]. Since $T(rL)$ is surjective, $T(rL)^{*}$ is bounded below, and thus the corona theorem guarantees the existence of a right inverse. ∎
We note that if $f$ is an NC entire function then the operator $f(rL)$ makes sense for every $r>0$. Moreover, since the power series of $f$ converges uniformly on any ball, we have $f(rL)\in\mathbb{A}_{d}$. Here, $\mathbb{A}_{d}$, the NC or free disk algebra of Popescu is the norm closure of the free algebra in $\mathbb{H}^{\infty}_{d}$. Equivalently, $\mathbb{A}_{d}$ is completely isometrically isomorphic to the unital norm closed algebra generated by the left creation operators in $\mathscr{B}(\mathbb{H}^{2}_{d})$. We will typically view elements of $\mathbb{A}_{d}$ and $\mathbb{H}^{\infty}_{d}$ as left multiplication operators on $\mathbb{H}^{2}_{d}$ and given any $F\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{m\times n}$ we will sometimes write $F(L)$ in place of $F$, to denote that $F$ acts as left multiplication by $F$. Indeed, evaluating the free formal power series of $F$ at $L=(L_{1},\cdots,L_{d})$ yields the operator of left multiplication by $F$ (if $F\in\mathbb{A}_{d}\otimes\mathbb{C}^{m\times n}$ the partial Cesàro sums of this power series converge in operator-norm, while if $F\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{m\times n}$ then these Cesàro sums and their adjoints converge in the strong-* operator topology, as was discussed previously).
@@
Let $F=(f_{1},\cdots,f_{n})\in\mathscr{O}_{d}(R)^{1\times n}$, $R\in(0,+\infty]$. For every $0<r<R$, let $F(rL)=V_{r}(L)T_{r}(L)$ be the inner–outer decomposition. Then, there exists $0<\rho<R$ and $\leq n, kk\in\mathbb{N}$, such that for every $r>\rho$, the length of $V_{r}$ is $k$. Moreover, if $\rho<r<s<R$, then $V_{s}(\tfrac{r}{s}L)$ is injective and $\mathrm{Ker}\,F(rL)=\mathrm{Ker}\,T_{s}(\tfrac{r}{s}L)$.
@@
Since the dimension of the wandering subspace is at most the number of components, we obtain that $w(r)\leq w(s)$. In other words, the function $w(t)$ is monotonically increasing, bounded, and integer–valued. Therefore, there exists a $\rho>0$, such that for all $0<\rho<r<R$, $w(r)$ is constant.
For $r>\rho$, we apply the inner–outer decomposition $V_{s}(\tfrac{r}{s}L)=V(L)H_{s,r}(L)$, where $H_{s,r}$ is outer and square. Observe that any square outer is necessarily injective. Indeed, at every point $Z\in\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, we have that $H_{s,r}(Z)$ is surjective and, therefore, invertible. We conclude that any vector in the kernel of $H_{s,r}$ must evaluate to 0 on $\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, and thus, is zero. Therefore, since $V(L)$ is an isometry, $V_{s}(\tfrac{r}{s}L)$ is injective.
The final statement then follows as if $V_{s}(\tfrac{r}{s}L)$ is injective, $T_{s}(\tfrac{r}{s}L)$ is surjective and $F(rL)=V_{s}(\tfrac{r}{s}L)T_{s}(\tfrac{r}{s}L)$ is injective, we must also have that $T_{s}(\tfrac{r}{s}L)$ is injective. By the open mapping theorem, $T_{s}(\tfrac{r}{s}L)$ is then invertible. Any invertible $A(L)\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{m\times n}$ must be square, i.e. $m=n$. Indeed, $A(L):\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{n}\rightarrow\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{m}$ is surjective. Hence $\{A(L)1\otimes e_{k}\}_{k=1}^{n}$ is a $R\otimes I_{m}$-cyclic set for $\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{m}$ and hence $m\leq n$. Conversely, $A(L)^{-1}:\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{m}\rightarrow\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{n}$ is also surjective, so $m=n$. ∎
@@
[25, Appendix] Let $E\in\mathbb{H}^{\infty}_{d}\otimes\mathbb{C}^{n\times n}$ be idempotent. Then there exists an $S\in\operatorname{GL}_{n}(\mathbb{H}^{\infty}_{d})$ and a $1\leq k\leq n$ so that
@@
This lemma was established in [25, Appendix], we include the proof for completeness.
@@
Write $S(L):=(V(L),W(L))$. Since $E(L)^{2}=E(L)$, we obtain $V(L)(I_{\mathbb{H}^{2}_{d}}\otimes I_{k}-F(L)V(L))F(L)=0$. Since $V(L)$ is injective and $F(L)$ has dense range, it follows that $F(L)V(L)=I_{\mathbb{H}^{2}_{d}}\otimes I_{k}$. Similarly, $G(L)W(L)=I_{\mathbb{H}^{2}_{d}}\otimes I_{j}$. Moreover, since E(L)(IE(L))=0=(IE(L))E(L)E(L)(I-E(L))=0=(I-E(L))E(L), we conclude $F(L)W(L)=G(L)V(L)=0$. Hence, $S(L)$ is invertible with inverse $S(L)^{-1}=\left(\begin{smallmatrix}F(L)\\ G(L)\end{smallmatrix}\right)$. In particular, both are square, i.e., $j+k=n$.
@@
One of our main results, Theorem 4.3, proves that each $\mathscr{O}_{d}(R)$, $R\in(0,+\infty]$, is a semifir, and hence is projective–free. That is, a ring is projective–free if every finitely generated projective module over the ring is a free module. This implies, in particular, that any idempotent $E\in\mathscr{O}_{d}(R)^{n\times n}$ has such a factorization. Indeed, given such an $E$, we have that $\mathrm{Ran}\,E$ and $\mathrm{Ran}\,I-E$ are both finitely–generated projective right $\mathscr{O}_{d}(R)$-modules contained in $\mathscr{O}_{d}(R)^{n}$ obeying $\mathscr{O}_{d}(R)^{n}=\mathrm{Ran}\,E+\mathrm{Ran}\,I-E$ and $\mathrm{Ran}\,E\cap\mathrm{Ran}\,I-E=\{0\}$. Since $\mathrm{Ran}\,E$, $\mathrm{Ran}\,I-E$ are fintely–generated and $\mathscr{O}_{d}(R)$ is semifir, these are both free right $\mathscr{O}_{d}(R)$-modules of finite rank. Choosing a basis $\{f_{1},\cdots,f_{k}\}\subseteq\mathscr{O}_{d}(R)$ for $\mathrm{Ran}\,E$ and a basis $\{f_{k+1},\cdots,f_{n}\}\subseteq\mathscr{O}_{d}(R)$ for $\mathrm{Ran}\,I-E$, it follows that $\{f_{i}\}_{i=1}^{n}$ is a basis for $\mathscr{O}_{d}(R)^{n}$. Defining the matrix $S\in\mathscr{O}_{d}(R)^{n\times n}$ by
@@
Let $S\in\operatorname{GL}_{n}(\mathbb{H}^{\infty}_{d})$ and let $\mathbb{H}^{2}_{d}\otimes\mathbb{C}^{n}=\mathcal{H}\dotplus\mathcal{K}$ be a decomposition into a non-orthogonal closed direct sum of right invariant subspaces. Then, $\dim\mathscr{W}_{\mathcal{H}}+\dim\mathscr{W}_{\mathcal{K}}=n$, $\dim\mathscr{W}_{S\mathcal{H}}=\dim\mathscr{W}_{\mathcal{H}}$, and $\dim\mathscr{W}_{S\mathcal{K}}=\dim\mathscr{W}_{\mathcal{K}}$.
[…]
@@
The first claim is proved in [25, Appendix] and in Lemma 2.6 above. For the second claim, we note that $S\mathscr{W}_{\mathcal{H}}$ generates $S\mathcal{H}$ by right shifts. Therefore, $\dim\mathscr{W}_{S\mathcal{H}}\leq\dim\mathscr{W}_{\mathcal{H}}$. Now applying this inequality to $S^{-1}$, we get the equality. Similarly, for $\mathcal{K}$. ∎
The following lemma is a variation of [25, Lemma 5.2].
@@
Let $0<r<s<R$. Then the map $\Phi_{r/s}$ will map $F(sL)$ to $F(rL)$. Therefore, for every $h\in\mathrm{Ker}\,F(sL)$, we have that $\Phi_{r/s}(h)\in\mathrm{Ker}\,F(rL)$. Hence, $\mathcal{H}:=\overline{\Phi_{r/s}(\mathrm{Ker}\,F(sL))}$ is a closed right-invariant subspace of $\mathrm{Ker}\,F(rL)$. In particular, $\mathcal{W}:=\Phi_{r/s}(\mathscr{W}_{\mathrm{Ker}\,F(sL)})$ is a generating set for $\mathcal{H}$. Let $P$ be the projection onto $\mathscr{W}_{\mathcal{H}}$. Let $u\in\mathscr{W}_{\mathcal{H}}\ominus P\mathcal{W}$. Fix an arbitrary $w\in\mathcal{W}$. Then,
@@
Since $\mathcal{H}\ominus\mathscr{W}_{\mathcal{H}}$ is $R$-invariant, the second summand on the left is 0. Since $R^{\alpha*}u=0$ for every non-empty $\alpha$ and $u\perp w$ by assumption, the first summand is also 0. Finally, since $\mathcal{W}$ is a generating set, $u=0$, and therefore $\mathscr{W}_{\mathcal{H}}=\overline{P\mathcal{W}}$. ∎
@@
In [25, Appendix], it is proved that $\mathbb{H}^{\infty}_{d}$ is a projective–free ring. However, it is also claimed incorrectly in [25, Appendix] that $\mathbb{H}^{\infty}_{d}$ is a semifir. Indeed, this fails even for $d=1$, as can be seen in [54, Beispiel 2]. Namely, if we take $f(z)=e^{\frac{z+1}{z-1}}$ which is a (singular) inner function and g(z)=1zg(z)=1-z, an outer function, then by the corona theorem $f$ and $g$ do not generate the trivial ideal. Moreover, the ideal generated by $f$ and $g$ is not principal. The reason is that if $h=h_{i}\cdot h_{o}$ is the inner–outer factorization of an $h$ that divides both $f$ and $g$, then we since $f=ha$, $h_{o}$ must be invertible since $f$ has no outer part. On the other hand, since $g=hb$, $h_{i}=1$ since $g$ has no non-trivial inner part. Hence, $h$ is a unit, but this contradicts the fact that the ideal generated by $f$ and $g$ is not trivial.
As we will see, the obstruction that prevents $\mathbb{H}^{\infty}_{d}$ from being a semifir is that it is generally not possible to “trivialize” row outers that have dense range in the sense of [10, Section 2.3 and Theorem 2.3.1]. In this paper, we will see that for outers this problem can be overcome by decreasing the radii of the balls. Namely, any such row outer on a row-ball of a given radius can always be trivialized on any ball of strictly smaller radius. Even though $\mathbb{H}^{\infty}_{d}$ is not a semifir, it would be interesting to determine whether $\mathbb{H}^{\infty}_{d}$ is a Sylvester domain, which is a larger class of rings than semifirs that also have universal skew fields of fractions [10, Section 5.5].